OpenStudy (anonymous):

A mountain has a rise of 8,480 feet from its base. If there is a 2 degree drop in temperature for every rise of 530 feet in altitude, how many degrees difference is there between the ground temperature and the temperature at the peak?

OpenStudy (anonymous):

\[\frac{8480}{530}2=32{}^{\circ} \]

OpenStudy (anonymous):

(8480 / 530) * 2 = 16 * 2 = 32 degrees